# Evaluating Grounded Reasoning by Code-Assisted LLMs for Mathematics

### v1

@@ 3.5 Automated Evaluation
Figure 2: Confusion matrices, showing counts, of Code-Structure judge (left) and LLM-judge (right) on the held-out set of the annotated data from ASDiv. Overall accuracies are $81\%$ and $73\%$ for Code-Structure judge (ours) and LLM-judge (OpenAI o3-mini), respectively. ’FS Imp’: From-scratch implementation, ’BF’: Brute Force.
Extending the human manual analysis to the entire evaluation dataset consumes a lot of time and effort. Therefore, we look into automating the evaluation by training classifiers that label generated programs using one class from our proposed taxonomy. For this purpose, we first annotate a training set using categories from the taxonomy, then we train a decision tree classifier using features extracted from the programs. We compare the trained classifier to an LLM-Judge and evaluate both judges on a held out annotated set. We pick the more accurate judge for the large-scale evaluation.
@@ 3.5.1 Training Set Annotation
To train the automated evaluation mehtods, two authors annotate a randomly sampled subset of the generated programs using labels from the taxonomy. The annotated set is $300$ programs from ASDiv split. $210$ examples were used for training, while the rest are held out for measuring the performance of the automated judges.
The annotation guidelines were developed based on observations from the manual analysis. After experimentation with different annotations schemes, we found that human assignment of per-program labels produced low inter-annotator agreement (IAA), while per-line labeling produced very high agreement. We thus selected a per-line annotation scheme, along with a simple algorithm for assigning a program level label based on the per-line labels. More on the annotation guidelines and annotators background can be found in Appendix B.
@@ 3.5.2 Automated Evaluation Methods
  • Number of variables defined but not used. Number of variables used but not defined.
@@
Following other related work that utilize LLMs as judge Tong and Zhang (2024), we employ an LLM to assess a given program and assign a taxonomy class to that program. For this purpose, we provide a detailed description of the taxonomy classes in the prompt, and ask the LLM-judge to carefully analyze the input program, finally, we ask it to provide a single class number that fit the program most. The prompt includes no reference programs of taxonomy classes. The full prompt is provided in Appendix A.4.
@@ 4.1 Evaluated Models
We analyze code generations of several LLMs, including open-source models such as: StarCoder2 15B Lozhkov et al. (2024), Llama 3.1 8B Dubey et al. (2024) and Qwen2.5 7B Qwen et al. (2025). We use the instruction-tuned version of these models. Furthermore, we evaluate GPT-4 Achiam et al. (2023) and GPT4o-mini from OpenAI. All models were evaluated with greedy decoding. Implementation details are in Appendix A.1
@@ 4.2 Prompts
We adept the prompt from Gao et al. (2023) to evaluate LLMs on the ASDiv dataset, with only three demonstrations rather than eight, as no further gain was observed with the full set of demonstrations. For the MATH500 data, we prompt LLMs using demonstrations from the training split of the dataset, adapted from Gou et al. (2023). We evaluate GPT4o-mini and Qwen in zero-shot settings, as we observe models tend to generate less generations in natural language using this setting. The two full prompts can be found in Appendix A.3.
Model Conceptual Primitive FS Implementation Brute-Force Disorganized No Logic
GPT4o-mini $90.5\%$ $96.5\%$ $86.3\%$ $77.7\%$ $93.7\%$ $85.7\%$
GPT4 $86\%$ $86\%$ $100\%$ $68.4\%$ $60\%$ $63.6\%$
Qwen2.5 $91.5\%$ $78.3\%$ $89.4\%$ $100\%$ $60\%$ $100\%$
StarCoder2 $48\%$ $44.4\%$ $76.4\%$ $47.6\%$ $30.7\%$ $61.5\%$
Llama3.1 $35.7\%$ $48.1\%$ $37.5\%$ $53.7\%$ $33.3\%$ $57.5\%$
Table 1: Execution (macro) accuracy in percentages per taxonomy class for all evaluated models on the ASDiv math problems. ’FS’ is From Scratch.
5 Results and Analysis
5.1 Comparison of different automated methods
@@
To compare automated evaluation methods, we measure the agreement to human judgment using accuracy to ground truth labels assigned by annotators to each generated program.
Code-Structure Judge achieved a mean accuracy of $81\%$ with a standard deviation of $0.005$ over four different seeds. Figure 2 (left) shows the confusion matrix of the Code-Structure judge on the held out set from ASDiv data.
@@
We notice that Code-Structure judge achieves both high precision and recall on most of the classes, except for the Disorganized class with a low recall of $0.16$. The Disorganized programs can be a mix of other classes and the distinction between these can be semantic rather than structural. Class Primitives also has a lower precision than other classes with $0.72$. Appendix C.2 include some incorrectly classified programs by the Code-Structure judge, and the full table of precision, recall and F1-score.
LLM-judge, on the other hand, achieved lower accuracy with only $73\%$. The Primitives class has a low precision of $0.56$, while the Disorganized and No Logic classes have much lower recall of $0.25$ and $0.37$ respectively, decreasing the overall accuracy of this judge by $8\%$ in comparison to the Code-Structure judge. Figure 2 (right) shows the confusion matrix of LLM-judge in comparison to ground truth labels.
We conducted a qualitative error analysis of some incorrectly classified instances of the No Logic class and found that the LLM-judge tends to mistake the step of transcribing information from the question to be part of the logic, and consequently classifies the instance as Primitive instead.
5.2 Insights on code-assisted LLMs from evaluating their intermediate programs
@@
We employ the Code-Structure judge for large-scale evaluation of the entire set of generated programs in response to math word problems, aiming to answer our research questions:
Do code-assisted LLMs ground their generations to math rules?
The distribution of taxonomy classes per model on ASDiv is presented in Figure 3. We observe that while all models generate mathematically-grounded programs, there are differences in the way they ground: GPT models and Qwen mainly ground through library calls, while the other open-source models ground through primitives. All models depend less on from-scratch implementation. Older models (StarCoder2, Llama3.1, and GPT4) generate more ungrounded programs (29%, 43%, and 26%, respectively) as compared with the newer models (GPT4o-mini 14% and Qwen 7%). The majority GPT4 and Llama3.1 ungrounded generations are “Brute-force search” (20% and 26% of all generated programs, respectively), while StarCoder2 produces more “No logic” solutions (16% of all generations). All models generate very small percentage of disorganized programs.
@@
Figure 3: Distribution of taxonomy classes for all evaluated models in percentages on ASDiv problems. Classes are: ’Con’: Conceptual, ’Prim’: Primitives, ’FS Imp’: From scratch Implementation, ’BF’: Brute-Force, ’D’: Disorganized, and ’NoL’: No Logic.
What is the relation between grounding generations to math rules and end performance?
Table 1 provide the execution macro accuracy per taxonomy class for all evaluated models on the ASDiv math problems. We note that GPT models and Qwen grounded programs achieve higher accuracy than the grounded programs generated by less capable open-source models, indicating that grounding is not as effective for these models. On the other hand, many models perform well on the ungrounded programs; both Llama3.1 and StarCoder2 achieve higher accuracy on No Logic than Conceptual programs. This implies that these models attempt to solve math problems by memorizing the answers without solving the underlying reasoning problems. Similarly, many of the models employ Brute-Force searches to find the final answer, without guiding the search with any mathematical rules. Studying the correlation between the execution accuracy and grounding provided no useful insights, as models seems to perform well on all categories. [^2] Given the small sample size of ungrounded generations of GPT models and Qwen, we can’t conclude much about their performance on these categories. Overall, we can see that grounding generations to math knowledge is more effective for GPT models and Qwen. Conversely, open-source models depend on other hacks to approach math problems, which might indicate their inability to perform a successful grounding.
Model Conceptual Primitive FS Implementation Brute-Force Disorganized No Logic
GPT4o-mini $56.5\%$ $59.2\%$ $48.7\%$ $48.6\%$ $64.2\%$ $56.8\%$
Qwen2.5 $48.6\%$ $54.2\%$ $54.5\%$ $60.2\%$ $23.5\%$ $62.5\%$
StarCoder2 $10\%$ $17.1\%$ $0.0\%$ $25.0\%$ $8.3\%$ $46.1\%$
Table 2: Execution (macro) accuracy in percentages per taxonomy class for three evaluated models on the MATH500 problems. ’FS’ is From Scratch.
Why do LLMs perform differently within a class?
GPT4o-mini has very high accuracy on the Disorganized class, unlike all other models. To further understand this counter-intuitive phenomenon, we analyze the $16$ generations in this class by GPT4o-mini and find that out of all generations, $14$ are false positives, with $9$ Primitives and $5$ No Logic. We also investigate the reason why Llama3.1 and StarCoder2 generally have degraded performance on the grounded classes. On the Conceptual class, we qualitatively analyze 20 programs by the two models and find that 18/20 programs invoke the wrong concept, indicating a poor ability to invoke relevant math concepts. The other 2/20 programs incorrectly transcribe information from the question. Furthermore, on the Primitive class, we conduct error analysis on 20 programs generated by the two models and find that 9/20 were false positives of Disorganized and No Logic, and 8/20 have inferred step error, where the model infer the calculation of one step rather than generating the steps to calculate it, causing the error to propagate through the rest of the code. Additionally, 2/20 incorrectly transcribed information from the question and 1/20 had wrong ground truth answer. Finally, on the class From-scratch Implementation, Llama3.1 has the lowest accuracy in comparison to other models. Examining 10 random generations from this class, we find that the model implements a different solution than what the problem requires. For example implementing the code for the Greatest Common Divisor problem, when the question is asking for the Least Common Multiple or vice versa. Starcoder2, on the other hand, successfully implements many solutions.
How do more challenging math problems affect LLMs’ grounding?
To answer this question, we evaluate programs generated by GPT4o-mini, Qwen2.5 and StarCoder2 in response to problems from the MATH500 dataset, which has been found to be more challenging for LLMs Hendrycks et al. (2021). Figure 4 illustrates the distribution of taxonomy classes per evaluated LLM on MATH500.
@@
Figure 4: Distribution of taxonomy classes for all evaluated models in percentages on MATH500 dataset. Classes are: ’Con’: Conceptual, ’Prim’: Primitives, ’FS Imp’: From scratch Implementation, ’BF’: Brute-Force, ’D’: Disorganized and ’NoL’: No Logic.
We notice that more challenging data changes the distribution of grounded generations. The number of conceptual programs decreased to half for all models, while the number of ungrounded programs, specifically Brute-Force and No Logic, doubled for all models. This indicates that the difficulty of the problem might push models to rely less on grounded program.
As the number of programs in the No Logic class increased, we qualitatively analyze some generations in this class and find that for StarCoder2 and Qwen2.5 many instances actually just return the answer right away, while for GPT-4o-mini, many instances solve the questions with comments in natural language, rather than programmatic steps that can be executed to get the answer. We still consider these instances to be No Logic as these instances resemble reasoning in natural language, and code is not efficiently helping in any way in finding the correct answer. The preference of textual reasoning over code might be due to problem complexity, as empirically investigated and discussed in (Chen et al., 2025), where they show that some GPT models prefer textual reasoning over code reasoning depending on the complexity of the task.
As for the relation between execution accuracy and grounding, Table 2 presents execution macro accuracy per class per model on the MATH500 problems. We observe that the accuracy on ungrounded programs is higher or on par with accuracy on grounded generations, which is more pronounced for both open-source models  [^3] The high accuracy on Disorganized class for GPT4o-mini and Qwen is due to false positives.. This is problematic, as relying on memorized information or brute-force searches to find the answer is not trustworthy, and can’t provide consistent performance. We also study the library calls made by the evaluated LLMs on the two datasets and find that GPT4o-mini and Qwen exploit a greater variety of library calls in both datasets, but the distribution of those library calls shifts from ASDiv (where math.gcd, abs, math.lcm, min are most frequent) to MATH500 (where sympy.solve dominates). Starcoder2, on the other hand, doesn’t employ many libraries and barely generates Conceptual programs on the MATH500 dataset.
Finally, we look at the relation between math subjects and taxonomy classes to check if models generate a certain type of programs depending on the subject of the given problem. Both GPT4o-mini and Qwen2.5 tend to use more Primitive solutions for easier problems, such as Prealgebra and Precalculus. Additionally, they employ Conceptual programs for Algebra and Probability problems. Notably, both models also exploit memorized information or brute-force search for Intermediate Algebra, Number Theory, and some of the Algebra problems. Figure 9 in Appendix D.2 illustrates the distribution of taxonomy classes per math subject from the MATH500 dataset. StarCoder2, on the other hand, heavily depends on the Primitive solutions for most of the subjects, demonstrating a lack of diversity in the logic it employs for different types of math problems.
@@ 6 Conclusion
In this work, we conducted an in-depth analysis of code-assisted LLMs generated programs in response to math problems. Our assessment focuses on the extent to which LLMs employ math concepts in the generated programs and how this affects their end performance. We categorized the generated programs into six different categories, which make our proposed taxonomy of grounding to math concepts. We annotated a corpus of programs using classes from the taxonomy and trained a decision-tree model that we call Code-Structure Judge. The Code-Structure judge outperformed an LLM-judge baseline and was employed for a large-scale evaluation of more generated programs. Our findings show that models employ grounded reasoning, as well as ungrounded reasoning, comprising the use of brute-force search and memorized information. The use of ungrounded hacks depends highly on LLMs capabilities and problems difficulty. Our work emphasizes the need for holistic evaluation of code-assisted LLMs, and we hope our findings will inspire future work to further study why LLMs employ ungrounded solutions and how to mitigate this phenomenon.
@@
We note there’s limitations to our work. First: Code-structure Judge is still not accurate on the Disorganized class, causing many false positives. Second, we depend on one prompting technique and don’t compare performance when utilizing prompts to improve generated programs with further refinement, such as Self-Refine (Madaan et al., 2024) or Code-based Self-Verification (Zhou et al., 2023a).

### v2

@@ 3.6 Automated Evaluation
Figure 3: Confusion matrices, showing counts, of Code-Structure judge (left) and LLM-judge (right) on the held-out set. Overall accuracies are $81\%$, with a standard deviation of $0.005$ over four different seeds, and $73\%$ for Code-Structure judge (ours) and LLM-judge (OpenAI o3-mini), respectively. ’FS Imp’: From-scratch implementation, ’BF’: Brute Force.
Extending human manual analysis to the entire evaluation dataset requires a significant amount of time and effort. Therefore, we investigate automating the evaluation by training classifiers that label generated programs using a single class from our proposed taxonomy. For this purpose, we first annotate a training set using classes from the taxonomy, then we train a decision tree classifier using features extracted from the programs. We compare the trained classifier to an LLM-Judge and evaluate both on a held-out annotated set. We pick the more accurate judge for the large-scale evaluation.
@@ 3.6.1 Training Set Annotation
To train the automated evaluation methods, two authors annotate a randomly sampled subset of the generated programs using labels from the taxonomy. The annotated set is $300$ programs, $210$ examples were used for training, while the rest are held out for measuring the performance of the automated judges. The annotation guidelines were developed based on observations from the manual analysis. After experimentation with different annotation schemes, we found that human assignment of per-program labels produced low inter-annotator agreement (IAA), while per-line labeling produced very high agreement. We thus selected a per-line annotation scheme, along with a simple algorithm for assigning a program-level label based on the per-line labels. More on the annotation guidelines, process, and annotators’ background can be found in Appendix B.
@@ 3.6.2 Automated Evaluation Methods
  • Number of variables defined but not used, and number of variables used but not defined.
@@
Following other related work that utilizes LLMs as a judge Tong and Zhang (2024), we employ an LLM to assess a given program and assign a taxonomy class to that program. For this purpose, we provide a detailed description of the taxonomy classes in the prompt and ask the LLM-judge to analyze the input program carefully. Finally, we ask it to provide a single class number that best fits the program. The prompt includes no reference programs for taxonomy classes. The full prompt is provided in Appendix A.4.
@@ 4.1 Evaluated Models
We analyze code generations of several LLMs, including open-source models such as: StarCoder2 15B Lozhkov et al. (2024), Llama 3.1 8B Dubey et al. (2024), and Qwen2.5 7B Qwen et al. (2025). We use the instruction-tuned version of these models. Furthermore, we evaluate GPT-4 Achiam et al. (2023) and GPT4o-mini from OpenAI. All models were evaluated with greedy decoding. Implementation details are in Appendix A.1
@@ 4.2 Prompts
We use the prompt from Gao et al. (2023) to evaluate LLMs on the ASDiv dataset, with only three demonstrations rather than eight, as no further gain was observed with the full set. For MATH500, we prompt LLMs using demonstrations from the training split of the dataset, adapted from Gou et al. (2023). We evaluate GPT4o-mini and Qwen in zero-shot settings, since these models generated more solutions in natural language otherwise. The two full prompts can be found in Appendix A.3.
5 Results and Discussion
5.1 Comparison of Automated Evaluation Methods
@@
To compare automated evaluation methods, we measure the agreement with human judgment using accuracy on the annotated held-out set.
@@
Code-Structure Judge achieved a mean accuracy of $81\%$. Figure 3 (left) shows the confusion matrix of the Code-Structure judge on the held-out. We notice that Code-Structure judge achieves both high precision and recall on most of the classes, except for the Disorganized class with a low recall of $0.16$. The Disorganized programs can be a mix of other classes, and the distinction between these can be semantic rather than structural. Appendix C.2 includes some incorrectly classified programs by the Code-Structure judge, and the table of precision, recall, and F1-score. LLM-judge, on the other hand, achieved lower accuracy with only $73\%$. The Primitive class has a low precision of $0.56$. In contrast, the Disorganized and No Logic classes have much lower recall of $0.25$ and $0.37$ respectively, decreasing the overall accuracy of this judge by $8\%$ in comparison to the Code-Structure judge. Figure 3 (right) shows the confusion matrix of LLM-judge in comparison to ground truth labels. We conducted a qualitative error analysis of some incorrectly classified instances of the No Logic class and found that the LLM-judge tends to mistake the step of transcribing information from the question to be part of the logic, and consequently classifies the instance as Primitive instead.
5.2 Evauation of Generated Programs of Code-assisted LLMs
@@
We employ Code-Structure judge for automatically evaluating the entire set of generated programs in response to various math problems, and provide the findings below:
LLMs’ capabilities significantly impact the type of implemented reasoning.
The majority of programs generated by GPT models and Qwen implement logically sound reasoning to solve input problems from the ASDiv dataset. These LLMs ground their programs in mathematical concepts, utilizing numerous API calls to math libraries. Additionally, for many questions, they employ only primitive operations, given the simplicity of some problems in the ASDiv dataset. Figure 4 illustrates this, where for the above-mentioned models, the Conceptual and Primitive classes are dominant in the distribution of all programs. On the other hand, the open-source models, StarCoder2 and Llama3.1, rely much more on Primitve programs but also on ungrounded, unsound reasoning hacks to implement the solution. For instance, Llama3.1 employs math libraries in only 11% of its programs, while resorting to Brute-Force searches and memorized information, demonstrated in No Logic class, in almost 40% of its generated programs.
@@
Figure 4: Distribution of programs logic for all evaluated models in percentages on ASDiv problems. Classes are: ’Con’: Conceptual, ’Prim’: Primitives, ’FS Imp’: From scratch Implementation, ’BF’: Brute-Force, ’D’: Disorganized, and ’NoL’: No Logic.
Complex math problems increase the generation of unsound reasoning for all evaluated LLMs.
The difficulty imposed by the MATH500 dataset completely alters the type of reasoning implemented in the generated programs. This dataset is reported to be challenging for LLMs Hendrycks et al. (2021), as it probs for skills such as Calculus, Geometry, Algebra, Probability, among others. GPT4o-mini and Qwen now generate 25% more programs with unsound, ungrounded reasoning, specifically, with exhaustive searches and programs lacking any logic, at the expense of Conceptual programs that utilize math libraries. Figure 5 demonstrates this phenomenon in the distribution of all generated programs. Upon qualitatively analyzing some programs generated by GPT-4o-mini in the No Logic class, we found that the drastic increase in the number of programs with this class can be attributed to the increase in programs with reasoning in the comments rather than code lines. We classify these instances as No Logic, because they resemble reasoning in natural language, and the code is not efficiently helping in any way to find the correct answer. The preference of textual reasoning over code might be due to problem complexity, as empirically investigated and discussed in (Chen et al., 2025), demonstrating that some GPT models prefer textual reasoning over code reasoning depending on the complexity of the task.
@@
Figure 5: Distribution of programs logic for all evaluated models in percentages on the MATH500 dataset. Classes are: ’Con’: Conceptual, ’Prim’: Primitives, ’FS Imp’: From scratch Implementation, ’BF’: Brute-Force, ’D’: Disorganized, and ’NoL’: No Logic.
5.3 Further Analyses
In this section, we study the impact of sound reasoning, or its absence, on LLMs’ end performance. Furthermore, we demonstrate how different math subdomains impact LLMs’ preferred logic for solving the problem.
Impact of type of implemented reasoning on LLMs End-performace.
To investigate the impact of the reasoning implemented in generated programs on LLMs’ end-performance on math tasks, we execute the generated programs and match the execution outcome to ground truth answers, then calculate execution accuracy over programs in each logic class. Table 1 presents execution accuracy per class on all datasets. On ASDiv, we observe that sound programs, from Conceptual, Primitive, and From-Scratch Implementation classes, score slightly higher accuracy than unsound programs. However, StarCoder2 and Llama3.1 fail to follow the same trend. Qualitative analysis of their generated programs indicates that while these two LLMs employ API calls or from-scratch implementation to solve the problems, they call or implement the wrong API functionalities, causing the observed low accuracy. On the challenging dataset MATH500, Table 1 illustrates that execution accuracy of programs is on par for all logic types, sound and unsound. This is problematic, as programs with logically unsound reasoning, i.e:, false positives, can’t be trusted or easily verified; instead, LLMs seem to hack their way to finding final answers. Finally, the reported high accuracy on the Disorganized programs is mainly due to some false positive predictions from the Code-Structure judge.
ASDiv
Model Conceptual Primitive FS Imp Brute-Force Disorganized No Logic
GPT4o-mini $90\%$ $96\%$ $86\%$ $77\%$ $93\%$ $85\%$
GPT4 $86\%$ $86\%$ $100\%$ $68\%$ $60\%$ $63\%$
Qwen2.5 $91\%$ $78\%$ $89\%$ $100\%$ $60\%$ $100\%$
StarCoder2 $48\%$ $44\%$ $76\%$ $47\%$ $30\%$ $61\%$
Llama3.1 $35\%$ $48\%$ $37\%$ $53\%$ $33\%$ $57\%$
MATH500
GPT4o-mini $54\%$ $57\%$ $41\%$ $46\%$ $64\%$ $51\%$
Qwen2.5 $37\%$ $59\%$ $51\%$ $52\%$ $23\%$ $56\%$
StarCoder2 $0\%$ $15\%$ $0\%$ $25\%$ $0\%$ $38\%$
Table 1: Execution (macro) accuracy in percentages per logic class for evaluated models on ASDiv (top) and MATH500 (bottom). Despite the unsound reasoning implemented in programs from Brute-Force and No Logic, high execution accuracy can still be achieved. High accuracy on Disorgnized programs is mainly due to false positive predictions from the Code-Structure judge. ’FS Imp’ is From-Scratch Implementation.
Impact of problem domain on type of implemented reasoning.
Given that MATH500 questions are annotated with the math subdomain they test for, such as Calculus, Algebra, Probability, etc, we investigate whether the evaluated LLMs consistently approach problems within a domain using the same logic. We observe that GPT4o-mini and Qwen2.5 tend to use more Primitive solutions for easier problems, such as Prealgebra. Additionally, they consistently employ Conceptual programs for Algebra problems. However, both models appear to be less consistent with the type of reasoning they employ across the rest of the subdomains, indicating higher uncertainty about the best logic to tackle the problems. Figure 6 illustrates the distribution of programs’ logic per MATH500 subdomains.
(a) GPT4o-mini
(b) Qwen2.5
(c) StarCoder2
Figure 6: Distribution of programs logic, showing counts, per math subdomain in MATH500. Classes: ’FS Imp’: From-Scratch Implementation, ’BF’: Brute-Force, ’Disorg’: Disorganized.
StarCoder2, on the other hand, heavily relies on Primitive solutions for all subdomains, demonstrating a lack of diversity in the logic it employs for different types of math problems.
@@ 6 Conclusion
In this work, we conducted an in-depth analysis of code-assisted LLMs generated programs in response to math problems. Our assessment focuses on evaluating the logical soundness of underlying reasoning processes implemented in the generated programs. We categorized the generated programs into six different categories, which make up our proposed taxonomy of logical soundness. Three of which represent sound reasoning that ground the programs in verifiable math concepts. In contrast, the other three exploit memorized information or exhaustive searches to find final answers. We annotated a subset of programs using classes from the taxonomy and trained a decision-tree classifier, which we call the Code-Structure Judge. The Code-Structure judge outperformed an LLM-judge baseline and was employed for a large-scale evaluation of more generated programs. Our findings show that the capabilities of LLMs and the difficulty of the problem impact the type of reasoning implemented, yet regardless, LLMs can exploit unsound reasoning to achieve comparable accuracy. Our work underscores the importance of a comprehensive evaluation of code-assisted LLMs. We hope our findings inspire future work to further study why LLMs employ ungrounded solutions and how to mitigate this phenomenon.
@@
We note there are limitations to our work. First: Code-structure Judge is still not accurate on the Disorganized class, causing many false positives. Second, we depend on one prompting technique and don’t compare performance when utilizing prompts to improve generated programs with further refinement, such as Self-Refine (Madaan et al., 2024) or Code-based Self-Verification (Zhou et al., 2023a).